# Open Your Ears to Take a Look: A State-of-the-Art Report on the Integration of Sonification and Visualization

### v1

@@ 3.4 Reading Levels
Rau et al. [$*{}^{*}$15] presented an audiovisual idiom that lets users interactively explore molecular structures by using a “virtual microphone” that can be placed inside a 3D molecular visualization. The visual design provides an overview using the “Whole” level while displaying details at group level via the sonification. At the same time, the design provides a user with information about potentially visually occluded data. Hence, it makes the user aware of the existence of a phenomenon, that they could, if relevant, study in detail at a later point. The metaphor of the microphone to be used to “listen into the data” is similar to the one of the stethoscope in [$*{}^{*}$22b], where the visualization provides an overview while users can interactively choose a subset of data to display acoustically. An example of an idiom working at the sonification “whole” level and the visualization “group’ level, is the Mode Explorer [YH17]. The design sonifies a high-dimensional data space while visualizing a two-dimensional projection of that space with a scatterplot.
Discussion: Out of all possible combinations of reading levels and display type, the majority of cases (27) use the “Whole” level for both the sonification and the visualization (some of them are [EEBR21, MMM18, RS22, SAR22, WW15]). The next biggest group in our database holds 24 papers that can display data at the “Whole” level using their visualization and at the “group” level using their sonification, such as [BF12, BTB23, FBC12, $*{}^{*}$18, $*{}^{*}$15]. Generally speaking, we can observe that the visualization parts are mostly covering the same or a higher level of reading than the sonification part of a design (compare table Table 3). This phenomenon can be directly related to the distribution of tasks on the two display types, with visualization more often providing an overview and the sonification rather displaying details. In section 6, we will argue for breaking such patterns in the future work of the community.
@@ 3.5 Search Level
Figure 4: The four search levels suggested by Munzner: lookup, browse, locate and explore. (Figure from “Visualization Analysis and Design” [Mun15] by Tamara Munzner, with illustrations by Eamonn Maguire, CC BY 4.0.)
@@
Inspired by the taxonomy suggested by Munzner [Mun15], we tagged four different search levels, again assigned individually for the sonification and the visualization parts of the idiom. The type of search a user applies depends on their prior knowledge. Users who know what they are looking for and where they can find it will do a lookup. Searching for an unknown pattern at a known location is called browsing while locating is a search without knowing the location but the pattern one is looking for. A search uninformed both with regards to the location and the type of pattern one is looking for is called exploration. Designs that combine sonification and visualization could serve each of the described search levels as well as a combination of them.
Malikova et al. [MAFP19b], for example, presented an idiom that helps users explore scalar fields using both their ears and their eyes. To do so, the regional magnitude of small areas of a scalar field is sonified using pitch. From the search level perspective, a user of this idiom does not generally know what to look or listen for at the beginning of the analysis, therefore both modalities are used for exploring the presented data. The system could, for example, reveal the existence of small symmetry differences in scalar fields. What visually seems to be a perfectly symmetrical field, could become apparent as not quite symmetric when two similar pitches result in an acoustic phenomenon called “frequency beating.” The clearly audible phenomenon makes a listener aware of the non-symmetry of the field, potentially resulting in them taking a closer look at their data.
Discussion: Generally speaking, many idioms can be used for more than one search level and it seems to be dependent on the prior knowledge of a user (both about the data and about the idiom) what search level they use it for. Nevertheless, the majority of papers in the surveyed literature offer the search level of “explore” and, in general, most papers use the same search level for both their visualization and their sonification parts (see Table 2 and Table 4). Some examples of designs employing the same search level are [CB17, $*{}^{*}$18, $*{}^{*}$22a, FBC12, Rib19, GR11, HK22], while fewer combine different search levels for the two modalities [HAR16, $*{}^{*}$15, RJ16, AK11, Bal15, CWM21].
@@ 3.6 Dataset Type and Level of Measurement
Table 6: The number of entries for the different level of measurement and their combinations between the two display techniques. Most idioms use identical level of measurement. Notably, 42 cases use a higher level of measurement with their sonification than with their visualization (numbers above the main diagonal). 20 cases use the levels of measurement the other way around (numbers under the main diagonal).
Two more categories we tagged individually for the sonification parts and the visualization parts are the dataset type (see Figure 5 and Table 2) and the level of measurement (see Table 2). We distinguish between the four dataset types of tables, networks, fields, and geometries [Mun15] and between the four levels of measurement of nominal, ordinal, interval, and ratio scale.
It is reasonable that most idioms display data from identical dataset types with both their visualization and their sonification. The most prominent dataset type is the table, used in designs such as [BTB23, DLVDCG22, $*{}^{*}$18, $*{}^{*}$22a, FBC12, HAR16, MMM18, RJ16]. Table 5 reveals a notable exception from the dominance of identical dataset types: the combination of visualized geometry in combination with a sonified tables, often related to maps and supplemented regional information [BF12, $*{}^{*}$23, MMU16, $*{}^{*}$12, RS22].
Concerning the levels of measurement of the displayed data, we observe, again, that most idioms display data from the same level (see Table 5). We list a selection of cases for ratio data  [$*{}^{*}$12, $*{}^{*}$18, $*{}^{*}$16, $*{}^{*}$22, Rön21], for interval data [BB19, DLVDCG22, Rib19, Her20, MMU16], for ordinal data [AK11, LF21, PCB23], and for nominal data [GKW21, HAR16, NSC16], and now will continue with a discussion of three methods to map the data to the senses, the levels of redundancy.
@@ 3.7 Level of Redundancy
Figure 6: The three levels of mapping-redundancy: Redundant mappings display the same information via visualization and sonification. Complementary mappings display different information to the two senses and mixed mappings map some information redundantly and other in a complementary way.
@@
Technical redundancy describes the technical mapping from the data to a visual and auditory representation. If all the displayed information is mapped to channels of both modalities, then the audiovisual display idiom employs a technically redundant mapping. An example is the auditory line graph: We see and hear identical information via spatial position (visualization) and pitch (sonification). In some cases, it is not enough to use a purely technical description of redundancy, without incorporating our way of perceiving the different displays as humans. Therefore, we introduce the term of "communicative redundancy."
Communicative redundancy describes the fact that technically redundant designs might encode the same information with different perceptual qualities. Hence, we could identify different patterns in data by listening to them, and then by looking at them. Communicative redundancy will usually be strongly related to technical redundancy, but there are exceptions such as the combination of a WAV form visualization (line chart) and an audification, such as in the Sonification of the Tohoku Earthquake in Japan in 2011, also described in [WW15]. Technically, those two are fully redundant (the same data attributes are displayed both visually and auditorily), but communicatively they complement each other.
@@
Dedicated environments that offer the design of audiovisual display idioms often employ redundant mappings [PC19, Rib18, CWM21, DLVDCG22, KLTW17, Rib19]. One of them is the Sonifigrapher (see Figure 3(c)), presented by García [Rib19], that combines a line chart and a parameter mapping sonification of the light curves from NASA’s exoplanet archive. The vertical position of a line is mapped to the pitch of a musical sound, essentially “playing back” the line from left to right (compare Figure 7).
@@
The work presented by Rau et al. [$*{}^{*}$15] complementarily enhances the visualization of molecular simulations by using auditory icons and parameter mapping sonification. A user can position a “virtual microphone” inside of a 3D rendered visual representation of a molecule and listen to processes that are visually occluded.
@@
Temor et al. [$*{}^{*}$21] presented a mixed mapping approach for the audiovisual analysis of computational fluid dynamics, in their specific case tailored towards the prediction of cerebral aneurysm ruptures, hence in the medical context. Their sonification design is psychoacoustically motivated in a way that amplifies the differences between different simulations of their spatiotemporally dynamic data. The authors explicitly mention their choice to apply a mixed mapping in their design. They motivate their decision by the observation that the ”spatiotemporal fluctuations are highlighted in a way that seems to be superior to the presentation of different information to different sensory modalities, which is in line with how we make sense of spatiotemporally-dynamic stimuli in everyday scenarios”. We consider this observation as highly plausible and inspirational for potential future research and will reflect upon it in the concluding discussion of this STAR.
@@
Maças & Martins and Machado present an audiovisual display idiom that displays consumption patterns collected from Portuguese supermarkets over the course of two years. The authors, just in line with Temor et al. [$*{}^{*}$21], also argue for the employment of a mixed design, their teams having made prior experience with complementary designs that seem to have been less effective.
@@
Discussion: Regarding the level of mapping-redundancy, a high-level observation is the following: When designing an audiovisual display idiom, the mixed mapping seems to show a special potential. The mapping overlap between the sonification and the visualization seems to help the user with perceiving an idiom as integrated, rather than as two displays existing next to each other. There are different options to design a mixed mapping, out of which a promising one seems to be to synchronize a visual animation with a sonification [$*{}^{*}$21, $*{}^{*}$18] . In similar manner synchronizing the spatial position of the visual display with the direction a user perceived the sound from can be helpful to perceptually integrate the two stimuli [CB17].
@@ 3.8 Evaluation Approaches
The evaluation of designs is a pressing issue in both the sonification and the visualization fields. Just as well, audiovisual display idioms need to be evaluated. Rönnberg and Forsell even argue for a future standardization of questionnaires that assess the usability of audiovisual representations that could be used in combination with other measurements [RF22]. To study the current practice of evaluation of audiovisual designs, we applied four of the classes of evaluation techniques suggested by Isenberg et al. [$*{}^{*}$13]. The four classes are user performance (UP), user experience (UE), algorithmic performance (AP), and qualitative result inspection (QRI).
@@
When evaluations objectively measure how specific features in a visualization or sonification affect user performance with a system, these approaches belong to the evaluation class UP [$*{}^{*}$13]. Controlled experiments using various time measurements and accuracy are typical example methods in this class. Rönnberg and Johansson [RJ16] explored the combination of visualization and sonification, where the users were tasked with identifying the visual area with the highest density in visual representations with and without support of sonification. The user performance was measured in accuracy and response time, and the results showed that the combination of visualization and sonification increased accuracy in comparison with visualization only but also that response times were longer when sonification was used.
@@
The class of UE includes evaluations that focus on subjective feedback and experiences on a visualization or sonification [$*{}^{*}$13]. The use of interviews and/or various questionnaires are common evaluation methods. In the corpus of papers this was a common approach for evaluating the audiovisual representations, and often used in combination with other evaluation approaches. The paper by Ballweg, et al., [BBV16] presents a study where users answered a questionnaire about their experience of an audiovisual system for drug design on a 5-point Likert scale. By comparing responses given before and after the study it could be determined how the system could be integrated into the users’ workflow, if sonification could have a positive effect on the task, and in what way the system could be improved.
@@
Few papers in the systematic literature review used AP as an evaluation approach. Evaluation approaches in this class should contain a quantitative study of the performance or quality of visualization and/or sonification algorithms [$*{}^{*}$13]. However, in the corpus of papers used in the classification, this approach was only found in a few cases and was employed to determine that an audiovisual algorithm could produce sufficient quality rather than determining a certain level of quality or comparing different algorithms. As an example, in the paper by Kariyado et al., [KAV21] it is stated that the objective evaluation showed that the system presented in the paper allowed for a default amount of 255 audio sources with any audio drop-outs. The assessment of the algorithm corresponding to a level of 255 was determined by the game engine Unity’s ability to produce 255 valid audio sources and not because of a systematic review of the number of audio sources.
@@
Evaluations in the class QRI aim to draw conclusions based on qualitative discussions and assessments of audiovisual representations [$*{}^{*}$13]. In contrast to UE, these types of evaluations do not involve end users but instead, ask the user of a visualization or sonification to assess the representation for themselves. In the work of Bru, et al., [BTB23], an approach to a combined audiovisual representation based on parallel coordinates or dense line charts is presented. In the paper several different attributes and characteristics in the representations are discussed based on the researchers’ own reflections but with no formal user evaluation performed.
@@
Discussion: As we have now demonstrated there are various approaches for evaluating audiovisual representations, and the aim of the work and the research questions asked to determine the class of evaluation chosen. In some cases, in the corpus of papers, two of these evaluation approaches have been combined to provide a better and more detailed analysis of the outcome of the study findings [BBV16, $*{}^{*}$16, $*{}^{*}$22]. Yet about a third of the studies included in the systematic review did not include an evaluation at all. The absence of evaluation might lead to a situation where promising audiovisual design ideas might be rejected as clear and convincing evidence of their usability, benefits, and function is not presented in a paper, or that less useful ideas are overrated and promoted.
@@ 3.9 Target Platforms and Interactivity
Audiovisual display idioms can be displayed on different target platforms and in different environments (see examples in Figure 7). A typical display for the combination of sonification and visualization is the computer desktop environment, as this is also the most commonly used environment for visualization-only designs. About two-thirds of the designs included in this survey are developed exclusively for a desktop environment.
@@
The second largest category of target platforms was physical and/or multi-user environments. This includes dome theaters and planetariums, which are commonly used for audiovisual display idioms toward topics of astronomy (as previously mentioned in subsection 3.1[HTHB22, RS22, EEBR21]. Another type of physical environment are dedicated immersive environments for collaborative data analysis, where several users can take part in the data exploration. One example of such an environment is the CRAIVE-Lab (Collaborative-Research Augmented Immersive Virtual Environment Laboratory), where a design visualized and sonified market data of 128 corporations by using the large panoramic display and the 128-channel loudspeaker array of the environment[CB17] (see Figure 7(b)).
@@
Figure 7: Examples of different target platforms and environments. (a) The combination of VR visualization and sonification for immersive exploration of noise in urban environments by Berger and Bill [BB19]. (b) An immersive virtual environment for audiovisual spatialized data sonification presented by Chabot and Braasch [CB17]. (c) A screenshot showing the interactive interface for sonified Hertzsprung-Russel diagram by Huppenkothen et al. [HPDW23]. All screenshots CC BY 4.0.
@@
A low-cost alternative for physical environments are virtual environments, which are most commonly enabled through head-mounted displays. An approach for an audiovisual display idiom in this type of environment is to let the user navigate the dataset in a first-person perspective and use spatialized sonification which dynamically changes depending on where the user is positioned in the dataset. An example is the design of Berger & Bill (see Figure 7(a)), which facilitates an immersive exploration of urban noise standards [BB19] by creating a virtual environment of a city where the sonification allows the user to listen to the collected noise levels by navigating the environment.
@@
The use of touch displays and other tangible interfaces for an audiovisual display idiom can allow for a unified integration of the sonification and the visualization. A recurring approach is to use the visualization as a graphical user interface which enables the sonification upon interaction. Through a touch display, the design by Ferguson et al. [FBC12] enables the user to filter a line graph with multi-touch gestures to select what data should be sonified. As a more analog approach, the system Sonophenology by Ness et al. [$*{}^{*}$12] lets the user select areas on a printed color-coded paper map through fiducial markers, which in turn selects what information should be conveyed by the sonification.
@@
About two thirds of the papers in the database include some form of interaction. One of the most recurring objectives of the interaction was to make a selection of the dataset that would be sonified [FBC12]. This can include selecting geographic regions which the dataset is attributed to [$*{}^{*}$12]. For example, the audiovisual display idiom of Matsubara et al. [MMU16] creates an interactive sonification system for exploration of seismic data through a horizontal and vertical range slider to specify a geographic region. Other forms of interaction include simpler tasks such as browsing sonifications of individual or groups of data objects [Rib19, $*{}^{*}$23]. Huppenkothen et al. [HPDW23] created an interactive multimedia version of the Hertzsprung-Russell Diagram, where a user can select individual data entries of the diagram to listen to their sonification. On the other hand, there also exist more complex forms of interactions such as navigating a 3D environment [BB19, JMP13, $*{}^{*}$18]. Papachristodoulou et al. [PBM15] created a design for navigating complex datasets of brain networks, where the sonification conveyed information of different brain regions of the dataset. Within the sonification community, it was the introduction of model-based sonification [Her02], that made user interaction an integral part of the data analysis process. An example of a model-based sonification is the Mode Explorer by Yang and Hermann [YH18], where scratching-interactions with a pencil on a scatter plot enables the user to investigate different modes and their properties through the sonification.
@@
Discussion: When displaying an audiovisual display idiom on a target platform other than a desktop environment, it often creates opportunities for interacting with the system in a novel manner [EEBR21, HK22, $*{}^{*}$12, PBM15]. In the context of data analysis, interaction is also a particularly relevant part of the user experience [BBV16, Bea11, EEBR21, HK22].

### v2

@@ 3.4 Reading Levels
Rau et al. [RFK15] presented an audiovisual idiom that lets users interactively explore molecular structures by using a “virtual microphone” that can be placed inside a 3D molecular visualization. The visual design provides an overview using the “whole” level while displaying details at the “group” level via the sonification. At the same time, the design provides a user with information about potentially visually occluded data. Hence, it makes the user aware of the existence of a phenomenon that they could, if relevant, study in detail at a later point. The metaphor of the microphone to be used to “listen into the data” is similar to the one of the stethoscope in [ERI22], where the visualization provides an overview while users can interactively choose a subset of data to display acoustically. An example of an idiom working at the sonification “whole” level and the visualization “group” level is the Mode Explorer [YH17]. The design sonifies a high-dimensional data space while visualizing a two-dimensional projection of that space with a scatterplot.
Discussion: Out of all possible combinations of reading levels and display type, the majority of cases (27) use the “Whole” level for both the sonification and the visualization (some of them are [EEBR21, MMM18, RS22, SAR22, WW15]). The next biggest group in our database holds 24 papers that can display data at the “Whole” level using their visualization and at the “group” level using their sonification, such as [BF12, BTB23, FBC12, GDAS18, RFK15]. Generally speaking, we can observe that the visualization parts are mostly covering the same or a higher level of reading than the sonification parts of a design (compare Table 3). This phenomenon can be directly related to the distribution of tasks on the two display types, with visualization more often providing an overview and the sonification rather displaying details. In section 6, we will argue for breaking such patterns in the future work of the community.
@@ 3.5 Search Level
Figure 4: The four search levels suggested by Munzner: lookup, browse, locate, and explore. (Figure from “Visualization Analysis and Design” [Mun15] by Tamara Munzner, with illustrations by Eamonn Maguire, CC BY 4.0.)
@@
Inspired by the taxonomy suggested by Munzner [Mun15], we tagged four different search levels, again assigned individually for the sonification and the visualization parts of the idiom. The type of search a user applies depends on their prior knowledge (see Figure 4). Users who know what they are looking for and where they can find it will do a lookup. Searching for an unknown pattern at a known location is called browsing while locating is a search without knowing the location but the pattern one is looking for. A search uninformed both with regards to the location and the type of pattern one is looking for is called exploration. Designs that combine sonification and visualization could serve each of the described search levels as well as a combination of them.
Malikova et al. [MAFP19b], for example, presented an idiom that helps users explore scalar fields. The regional magnitude of small areas of a scalar field is sonified using pitch. From the search level perspective, a user of this idiom does not generally know what to look or listen for at the beginning of the analysis, and therefore, both modalities are used to explore the presented data. The system could, for example, reveal the existence of small symmetry differences in scalar fields. What visually seems to be a perfectly symmetrical field could become apparent as not quite symmetric when two similar pitches result in an acoustic phenomenon called “frequency beating.” The clearly audible phenomenon makes a listener aware of the non-symmetry of the field, potentially resulting in them taking a closer look at their data.
Discussion: Generally speaking, many idioms can be used for more than one search level, and which one they are used for seems to be dependent on the user’s prior knowledge (both about the data and about the idiom). Nevertheless, the majority of papers in the surveyed literature offer the search level of “explore” and, in general, most papers use the same search level for both their visualization and their sonification parts (see Table 2 and Table 4). Some examples of designs employing the same search level are [CB17, DCM18, ERI22, FBC12, Rib19, GR11, HK22], while fewer combine different search levels for the two modalities [HAR16, RFK15, RJ16, AK11, Bal15, CWM21].
@@ 3.6 Dataset Type and Level of Measurement
Table 6: The number of entries for the different levels of measurement and their combinations between the two display techniques. Most idioms use identical levels of measurement. Notably, 42 cases use a higher level of measurement with their sonification than with their visualization (numbers above the main diagonal). Twenty cases use the levels of measurement the other way around (numbers under the main diagonal).
Two more categories that we tagged individually for the sonification parts and the visualization parts are the dataset type (see Figure 5 and Table 2) and the level of measurement (see Table 2). We distinguish between the four dataset types of tables, networks, fields, and geometries [Mun15] and between the four levels of measurement of nominal, ordinal, interval, and ratio scale.
It is reasonable that most idioms display data from identical dataset types with both their visualization and their sonification. The most prominent dataset type is the table, used in designs such as [BTB23, DLVDCG22, DCM18, ERI22, FBC12, HAR16, MMM18, RJ16]. Table 5 reveals a notable exception from the dominance of identical dataset types: the combination of visualized geometry in combination with a sonified tables, often related to maps and supplemented regional information [BF12, LSB23, MMU16, NRL12, RS22].
Concerning the levels of measurement of the displayed data, we observe, again, that most idioms display data from the same level (see Table 5). We list a selection of cases for ratio data  [ASH12, DCM18, GRK16, PFH22, Rön21], for interval data [BB19, DLVDCG22, Rib19, Her20, MMU16], for ordinal data [AK11, LF21, PCB23], and for nominal data [GKW21, HAR16, NSC16], and now will continue with a discussion of three methods to map the data to the senses, the levels of redundancy.
@@ 3.7 Level of Redundancy
Figure 6: The three levels of mapping-redundancy: Redundant mappings display the same information via visualization and sonification. Complementary mappings display different information to the two senses, and mixed mappings map some information redundantly and others in a complementary way.
@@
Technical redundancy describes the technical mapping from the data to a visual and auditory representation. If all the displayed information is mapped to channels of both modalities, then the audiovisual display idiom employs a technically redundant mapping. An example is the auditory line graph: We see and hear identical information via spatial position (visualization) and pitch (sonification). In some cases, it is not enough to use a purely technical description of redundancy without incorporating our way of perceiving the different displays as humans. Therefore, we introduce the term "communicative redundancy."
Communicative redundancy describes the fact that technically redundant designs might encode the same information with different perceptual qualities. Hence, we could identify different patterns in data by listening to them than by looking at them. Communicative redundancy will usually be strongly related to technical redundancy, but there are exceptions, such as the combination of a WAV form visualization (line chart) and an audification, such as in the Sonification of the Tohoku Earthquake in Japan in 2011, also described in [WW15]. Technically, those two are fully redundant (the same data attributes are displayed both visually and auditorily), but communicatively, they complement each other.
@@
Dedicated environments that offer the design of audiovisual display idioms frequently employ redundant mappings [PC19, Rib18, CWM21, DLVDCG22, KLTW17, Rib19]. One of them is the Sonifigrapher (see Figure 3(c)), presented by García [Rib19], that combines a line chart and a parameter mapping sonification of the light curves from NASA’s exoplanet archive. The vertical position of a line is mapped to the pitch of a musical sound, essentially “playing back” the line from left to right (compare Figure 7).
@@
The work presented by Rau et al. [RFK15] complementarily enhances the visualization of molecular simulations by using auditory icons and parameter mapping sonification. A user can position a “virtual microphone” inside of a 3D-rendered visual representation of a molecule and listen to visually occluded processes.
@@
Temor et al. [TMN21] presented a mixed mapping approach for the audiovisual analysis of computational fluid dynamics tailored towards the prediction of cerebral aneurysm ruptures, hence in the medical context. Their sonification design is psychoacoustically motivated in a way that amplifies the differences between different simulations of their spatiotemporally dynamic data. The authors explicitly mention their choice to apply a mixed mapping in their design. They motivate their decision by the observation that the ”spatiotemporal fluctuations are highlighted in a way that seems to be superior to the presentation of different information to different sensory modalities, which is in line with how we make sense of spatiotemporally-dynamic stimuli in everyday scenarios.” We consider this observation highly plausible and inspirational for potential future research and will reflect upon it in the concluding discussion of this STAR.
@@
Maças & Martins, and Machado present an audiovisual display idiom that displays consumption patterns collected from Portuguese supermarkets over the course of two years. The authors, just in line with Temor et al. [TMN21], also argue for the employment of a mixed design, their teams having made prior experience with complementary designs that seem to have been less effective.
@@
Discussion: Regarding the level of mapping redundancy, a high-level observation is the following: When designing an audiovisual display idiom, mixed mapping seems to show a special potential. The mapping overlap between the sonification and the visualization seems to help the user perceive an idiom as integrated rather than as two displays existing next to each other. There are different options to design a mixed mapping, out of which a promising one seems to be to synchronize a visual animation with a sonification [TMN21, MNW18]. In a similar manner, synchronizing the spatial position of the visual display with the direction that a user perceives the sound from can be helpful to perceptually integrate the two stimuli [CB17]. An audiovisual analysis idiom that utilizes complementary or mixed mappings enables a designer to choose the employment of sound instead of a second view. Data that would conventionally be made visible via a second view can be represented using sound instead, which can modify the design of the visual view itself. Such situations can become especially apparent whenever the visualization needs to fit on small screens such as on smartwatches.
@@ 3.8 Evaluation Approaches
The evaluation of designs is a pressing issue in both the sonification and visualization fields. Audiovisual idioms likewise need to be evaluated. Rönnberg and Forsell even argue for the standardization of questionnaires that assess the usability of audiovisual representations that could be used in combination with other measurements [RF22]. To study the current practice of evaluation of audiovisual designs, we applied four of the classes of evaluation techniques suggested by Isenberg et al. [IIC13]. The four classes are user performance (UP), user experience (UE), algorithmic performance (AP), and qualitative result inspection (QRI).
@@
When evaluations measure how specific features in a visualization or a sonification affect user performance with a system, these approaches belong to the evaluation class UP [IIC13]. Controlled experiments using various time measurements and accuracy are typical example methods in this class. Rönnberg and Johansson [RJ16] explored the combination of visualization and sonification, where the users had to identify the visual area with the highest density in visual representations with and without the support of sonification. The authors measured accuracy and response time, and the results showed that the combination of visualization and sonification increased accuracy in comparison with visualization only but also that response times were longer when sonification was used.
@@
The class of UE includes evaluations that focus on subjective feedback on and experiences with a visualization or a sonification [IIC13]. Interviews and/or various questionnaires are common evaluation methods. In our corpus of papers, this was a common approach for evaluating the audiovisual representations and was often used in combination with other evaluation approaches. The paper by Ballweg et al., [BBV16] presents a study where users answered a questionnaire about their experience of an audiovisual system for drug design on a 5-point Likert scale. By comparing responses given before and after the study, it could be determined how the system could be integrated into the users’ workflow, if sonification could have a positive effect on solving the task, and in what way the system could be improved.
@@
Few papers in the systematic literature review used AP as an evaluation approach. Evaluation approaches in this class should contain a quantitative study of the performance or quality of visualization and/or sonification algorithms [IIC13]. However, in our corpus of papers, this approach was only found in a few cases and was employed to determine that an audiovisual algorithm could produce sufficient quality rather than determining a certain level of quality or comparing different algorithms. As an example, in the paper by Kariyado et al. [KAV21], it is stated that the evaluation showed that the system presented in the paper allowed for a default amount of 255 audio sources with any audio drop-outs.
@@
Evaluations in the class QRI aim to draw conclusions based on qualitative discussions and assessments of audiovisual representations [IIC13]. In contrast to UE, these types of evaluations do not involve end users but instead, ask the user to assess the representation for themselves. In the work of Bru et al. [BTB23], an approach to a combined audiovisual representation based on parallel coordinates or dense line charts is presented. Several different attributes and characteristics in the representations are discussed based on the researchers’ own reflections, but no formal user evaluation has been performed.
@@
Discussion: We have now demonstrated there are various approaches for evaluating audiovisual representations. The aim of the work and the research questions that are asked determine the class of evaluation chosen. In some cases, in the corpus of papers, two of these evaluation approaches have been combined to provide a better and more detailed analysis of the outcome of the study findings [BBV16, GRK16, PFH22]. Yet about a third of the studies included in the systematic review did not include an evaluation at all. The absence of evaluation might lead to a situation where promising audiovisual design ideas might be rejected because clear and convincing evidence of their usability, benefits, and function is not presented in a paper, or less useful ideas are overrated and promoted.
@@ 3.9 Target Platforms and Interactivity
Audiovisual display idioms can be displayed on different target platforms and in different environments (see examples in Figure 7). A typical display for the combination of sonification and visualization is the computer desktop environment, as this is also the most commonly used environment for visualization-only designs. About two-thirds of the designs included in this survey are developed for a desktop environment.
@@
The second largest category of target platforms was physical and/or multi-user environments. This includes dome theaters and planetariums, which are commonly used for audiovisual display idioms toward topics of astronomy (as previously mentioned in subsection 3.1[HTHB22, RS22, EEBR21]. Another type of physical environment is a dedicated immersive environment for collaborative data analysis, where several users can take part in the data exploration. One example of such an environment is the CRAIVE-Lab (Collaborative-Research Augmented Immersive Virtual Environment Laboratory), where a design visualized and sonified market data of 128 corporations by using the large panoramic display and the 128-channel loudspeaker array of the environment[CB17] (see Figure 7(b)).
@@
Figure 7: Examples of different target platforms and environments. (a) The combination of VR visualization and sonification for an immersive exploration of noise in urban environments by Berger and Bill [BB19]. (b) An immersive virtual environment for audiovisual spatialized data sonification presented by Chabot and Braasch [CB17]. (c) A screenshot showing the interactive interface for the sonified Hertzsprung-Russel diagram by Huppenkothen et al. [HPDW23]. All screenshots CC BY 4.0.
@@
A low-cost alternative for physical environments is a virtual one, which is most commonly enabled through head-mounted displays. An approach for an audiovisual display idiom in this type of environment is to let the user navigate the dataset from a first-person perspective and use spatialized sonification, which dynamically changes depending on where the user is positioned in the dataset. An example is the design of Berger & Bill (see Figure 7(a)), which facilitates an immersive exploration of urban noise standards [BB19] by creating a virtual environment of a city where the sonification allows the user to listen to the collected noise levels by navigating the environment.
@@
The use of touch displays and other tangible interfaces for an audiovisual display idiom can allow for a unified integration of sonification and visualization. A recurring approach is to use the visualization as a graphical user interface that enables the sonification upon interaction. Through a touch display, the design by Ferguson et al. [FBC12] enables the user to filter a line graph with multi-touch gestures to select what data should be sonified. As a more analog approach, the system Sonophenology by Ness et al. [NRL12] lets the user select areas on a printed color-coded paper map through fiducial markers, which in turn selects what information should be conveyed by the sonification.
@@
About two-thirds of the papers in the database include some form of interaction. One of the most recurring objectives of the interaction was to make a selection of the dataset that would be sonified [FBC12]. This can include selecting a geographic region to which the dataset is attributed to [NRL12]. For example, the audiovisual display idiom of Matsubara et al. [MMU16] creates an interactive sonification system for exploration of seismic data through a horizontal and vertical range slider to specify a geographic region. Other forms of interaction include simpler tasks such as browsing sonifications of individual or groups of data objects [Rib19, LSB23]. Huppenkothen et al. [HPDW23] created an interactive multimedia version of the Hertzsprung-Russell Diagram, where a user can select individual data entries of the diagram to listen to their sonification. On the other hand, there also exist more complex forms of interactions such as navigating a 3D environment [BB19, JMP13, MNW18].
Papachristodoulou et al. [PBM15] created a design for navigating complex datasets of brain networks, where the sonification conveyed information about different brain regions of the dataset. Within the sonification community, it was the introduction of model-based sonification [Her02], that made user interaction an integral part of the data analysis process. An example of a model-based sonification is the Mode Explorer by Yang and Hermann [YH18], where scratching-interactions with a pencil on a scatter plot enable the user to investigate different modes and their properties through the sonification.
@@
Discussion: When displaying an audiovisual display idiom on a target platform other than a desktop environment, it often creates opportunities for interacting with the system in a novel manner [EEBR21, HK22, NRL12, PBM15]. In the context of data analysis, interaction is also a particularly relevant part of the user experience [BBV16, Bea11, EEBR21, HK22].